OpenStudy (anonymous):

Which of the following is the equation of a circle whose center is at the origin and whose radius is 5 ? A.) x(squared) + y(squared) =√5 B.) x(squared) + y(squared) = 5 C.) x(squared) + y(squared) = 10 D.) x(squared) + y(squared) = 25

OpenStudy (anonymous):

Are you familiar with the standard form equation for a circle?

OpenStudy (anonymous):

\[(x-a)^2+(y-b)^2=r^2\]

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

Where (a, b) is the coordinate of the circle's centre, and r is the radius

OpenStudy (anonymous):

your going to have to help me with it

OpenStudy (anonymous):

Well, you just need to plug in the information given from the problem.

OpenStudy (anonymous):

The circle is centered at the origin so \[(a,b) = (0,0)\]

OpenStudy (anonymous):

The radius is 5.

OpenStudy (anonymous):

\[(x -0)^2 + (y-0)^2 = 5^2\]

OpenStudy (anonymous):

how do i find x, and y?

OpenStudy (anonymous):

You're not trying to find x and y. You are asked to provide the equation of the circle in terms of x and y.

OpenStudy (anonymous):

So it would be D?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

i accendtly typed the radius wrong its supposed to be √5. would it still be 25?

OpenStudy (anonymous):

@math1234